In a certain region of space the potential is given by – V=k[2x2-y2+z2]. The electric field at the point (1,1,1) has magnitude.?

Grade:12

5 Answers

Vikas TU
10 years ago

E = - dV/dx

same with respecto y and z

now differentiate line wise and get the ansr.

Plz Approve!

Anmol Dhingra
10 years ago

aNS- 2 AND DIRN Z AXIS 

SOURAV MISHRA
10 years ago

use partial differential to determine the components of the electric field strength vector.

the electric field strength vector is given by 

E = -k(4xi^ - 2yj^ + 2zk^)

therefore at the point (1, 1, 1) the magnitude of the electric field strength is (24)(1/2).

Abhishek Parihar
10 years ago

the ans is-

[(4)^2 +(2)^2+(2)^2]^1\2

i.e.

(24)^1/2

Vishakha chandan
5 years ago
magnt. of E:  k(2-1+1)  / under root of (2^2 +(-1)^2 +1^2) =2k/under root of 6......Explain: put1,1,1 in eqn ..upon... put coefficient of x^2,Y^2,Z^2 in denomenator with eqn ...